# criterion and density

### v1

@@ 1.1. Main results
Our first contribution is a one-page proof, relying only on Dedekind’s index theorem, which determines $\alpha$-monogeneity in every pure degrees.
Criterion for $\alpha$-monogeneity. Let $K=\mathbb{Q}(\alpha)$ with $\alpha^{n}=m$, $n\geq 2$, $m\in\mathbb{Z}\setminus\{0\}$. Then
@@
Our second contribution is a clean distribution theory for $\alpha$-monogenic pure fields. Let $\mathcal{E}_{p}=\{m\in\mathbb{Z}:\ m^{p}\equiv m\pmod{p^{2}}\}$. By Hensel, among units modulo $p^{2}$ the “bad” classes are exactly the (p1)(p-1) Teichmüller lifts. Conditioning on square-freeness, a proportion $1/(p+1)$ of $m$ lie in $\mathcal{E}_{p}$, and the events $\{m\in\mathcal{E}_{p}\}$ are independent across primes dividing $n$ (Chinese Remainder + uniform square-free distribution in classes). Consequently, for the set of $m$ with $x^{n}-m$ irreducible (a density-$1$ restriction by Eisenstein), the natural density of $\alpha$-monogeneity is
@@
a product depending only on $\mathrm{rad}(n)$. We refine this density in arithmetic progressions (with explicit local factors), exhibit progressions on which $100\%$ of square-free $m$ are $\alpha$-monogenic, and count $\alpha$-monogenic pure fields by discriminant:
@@ 1.2. Related results
There is an extensive literature on monogeneity, integral bases, and pure fields, among others, see Gaál’s monograph [Gaá19] for a comprehensive treatment and his recent survey [Gaá24] for an updated overview. More generally, for radical extensions $L(\sqrt[n]{\alpha})/L$, Smith proved a relative criterion for $\sqrt[n]{\alpha}$ to generate a power integral basis. When specialized to $L=\mathbb{Q}$ and the pure family $x^{n}-m$, this yields exactly the two local conditions in Theorem 2.6 (square-freeness of $m$ and $\nu_{p}(m^{p}-m)=1$ for all $p\mid n$) [Smi21]. Our proof has the advantage of being uniform and short in the pure case (one pp‑adic evaluation in $(\mathbb{Z}/p^{2}\mathbb{Z})[X]/(X^{s}-m)$). See also [Gas17, EF21] for earlier formulations in the case $x^{p^{r}}-a$ and the associated Wieferich‑type obstruction $a^{p}\equiv a\pmod{p^{2}}$.
A companion paper by the first-named author studies the full $p$-adic shape of explicit integral bases for pure fields and proves that, for each $p^{e}\parallel n$, the local shape is determined by $a\bmod p^{e+1}$, yielding a minimal global period $M(n)=\prod_{p^{e}\parallel n}p^{e+1}=n\cdot\mathrm{rad}(n)$. In that framework the same local invariant $r_{p}(a)=\nu_{p}(a^{p-1}-1)-1$ governs denominators and gives a Wieferich-type threshold at $p\mid n$, which is exactly the quantity appearing in our $\alpha$-monogeneity criterion $\nu_{p}(m^{p}-m)=1$. See [ND25] for details.
The density of square-free integers in progressions and basic $p$-adic features of Frobenius (Teichmüller lifts) will play a role in our distributional results, for accessible references see Brown [Bro21] and standard sources such as Cohen [Coh93] or Koblitz [Kob84]. For pure cubics, monogeneity (not necessarily via $\alpha$) and quantitative aspects were studied by Aygin–Nguyen [AN21]. In the monic coefficient space, Bhargava–Shankar–Wang [BSW22, BSW25] proved that a positive density $\zeta(2)^{-1}$ of polynomials (f) satisfy $\mathbb{Z}[x]/(f)=\mathcal{O}_{\mathbb{Q}[x]/(f)}$, and they determined the density of squarefree discriminants. They further extend these results to binary forms and resolve the arithmetic Bertini case of $\mathbb{P}^{1}_{\mathbb{Z}}$. Our work addresses instead the thin family $f_{m}(x)=x^{n}-m$: we give a complete local–global criterion for $\mathbb{Z}[\sqrt[n]{m}]$ to be maximal and compute the exact density $\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}$ along the m-line. Thus, BSW treat the generic coefficient space (via invariant‑theoretic methods), while we handle a rigid one‑parameter family (via Dedekind’s criterion), and the two sets of results are complementary. From a moduli‑theoretic perspective, schemes parameterizing monogenic generators and notions of local monogenicity are developed in [ABS23], while not used here, this situates our local conditions in a geometric framework.
@@ 1.4. Acknowledgement
The first author thanks the Morningside Center of Mathematics, Chinese Academy of Sciences, for its support and a stimulating research environment.
@@ 2. A criterion for $\alpha$-monogenic pure fields
For example, when $K=\mathbb{Q}(\alpha)$ with $\alpha\in\{\sqrt{-2},\sqrt[3]{3},\sqrt[5]{6}\}$, one checks that $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$. We record a version of Dedekind’s index theorem as follows.
@@
If $n=p^{r}$ with $p$ prime and $r\geq 1$, then $K$ is $\alpha$-monogenic if and only if $m$ is square-free and $\nu_{p}(m^{p}-m)=1$.
Corollary 2.9.
@@ 3. Natural density for $\alpha$-monogenic pure fields
In this section we compute the natural density of integers $m$ for which $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ when $\alpha^{n}=m$ is pure. Throughout, $n\geq 2$ is fixed.
@@
A set $E\subset\mathbb{Z}$ has natural density $\delta(E)$ if the limit
$\delta(E):=\lim_{X\to\infty}\frac{1}{X}\#\{m\in E:\ 1\leq m\leq X\}$
exists (equivalently, the two-sided limit with $|m|\leq X$ exists and has the same value for the congruence- and squarefreeness-defined sets we consider). We write $\mu$ for the Möbius function, and $\mu^{2}$ for the indicator of square-free integers.
We will also use the following elementary observations.
@@
Let $g(x)=x^{p}-x$. In $\mathbb{F}_{p}$ the roots are all $x\in\mathbb{F}_{p}$, at each root mod $p$ one has $g^{\prime}(x)=px^{p-1}-1\equiv-1\pmod{p}$, which is invertible. By Hensel, each root lifts uniquely to $p^{k}$ for all $k$, so $\{x\bmod p^{2}:\ g(x)\equiv 0\}=p$. ∎
@@ 3.2. Global density via independence across primes
This is the key combinatorial identity underlying the main density theorem below.
We also record that the restriction $x^{n}-m$ irreducible is asymptotically harmless as follows.
@@
Contrast with the coefficient‑box model. Over the full coefficient space, Bhargava–Shankar–Wang prove that a positive proportion $\zeta(2)^{-1}$ of monic polynomials (or binary nn‑ic forms) define maximal orders [BSW22, BSW25]. In our family $f_{m}(x)=x^{n}-m$, Theorem 3.6 gives the exact maximal‑order density $\delta_{n}=(6/\pi^{2})\prod_{p\mid n}\frac{p}{p+1}$.
@@
Related density in thin families. For a different thin one‑parameter family, Bardestani computes the density of primes $p$ for which $t^{q}-p$ is monogenic, with complementary phenomena in the cubic case [Bar14].
@@ 3.3. Consequences and applications of the density formula
For questions beyond $\alpha$-monogeneity (e.g. tracking the full $p$-denominator profile of an integral basis), the canonical modulus is $M(n)=\prod_{p^{e}\parallel n}p^{e+1}=n\cdot\operatorname{rad}(n)$: the local $p$-shape depends only on $a\bmod p^{e+1}$ and this modulus is minimal. Our $\alpha$-monogeneity constraints at $p\mid n$ depend only on $a\bmod p^{2}$, but working modulo $M(n)$ aligns directly with the periodicity tables of the full shape, see [ND25].

### v2

@@ 1.1. Main results
Our first contribution is a one-page proof, relying only on Dedekind’s index theorem, which determines $\alpha$-monogeneity in every pure degrees (see Theorem 2.7). More precisely, let $K=\mathbb{Q}(\alpha)$ with $\alpha^{n}=m$, $n\geq 2$, $m\in\mathbb{Z}\setminus\{0\}$. Then
@@
Our second contribution is a clean distribution theory for $\alpha$-monogenic pure fields (see Theorem 3.6). Let $\mathcal{E}_{p}=\{m\in\mathbb{Z}:\ m^{p}\equiv m\pmod{p^{2}}\}$. By Hensel, among units modulo $p^{2}$ the “bad” classes are exactly the (p1)(p-1) Teichmüller lifts. Conditioning on square-freeness, a proportion $1/(p+1)$ of $m$ lie in $\mathcal{E}_{p}$, and the events $\{m\in\mathcal{E}_{p}\}$ are independent across primes dividing $n$ (Chinese Remainder theorem and uniform square-free distribution in classes). Consequently, for the set of $m$ with $x^{n}-m$ irreducible, the natural density of $\alpha$-monogeneity is
@@
a product depending only on $\mathrm{rad}(n)$. We refine this density in arithmetic progressions (with explicit local factors), exhibit progressions on which $100\%$ of square-free $m$ are $\alpha$-monogenic, and count $\alpha$-monogenic pure fields by discriminant (see Theorem 3.14)
@@ 1.2. Related results
There is an extensive literature on monogeneity, integral bases, and pure fields. For comprehensive treatments, we refer to the monographs of Evertse–Győry [EG17] and Gaál [GAÁ19]. See also the recent surveys by Gaál [GAÁ24] and Evertse–Győry [EG24] for updated accounts of current developments.
More generally, for radical extensions $L(\sqrt[n]{\alpha})/L$, Smith in [SMI21] proved a relative criterion for $\sqrt[n]{\alpha}$ to generate a power integral basis. When specialized to $L=\mathbb{Q}$ and the pure family $x^{n}-m$, this yields exactly the two local conditions in Theorem 2.7 (square-freeness of $m$ and $\nu_{p}(m^{p}-m)=1$ for all $p\mid n$). Our proof has the advantage of being uniform and short in the pure case (one pp‑adic evaluation in $(\mathbb{Z}/p^{2}\mathbb{Z})[X]/(X^{s}-m)$). See also [GAS17, EL 21] for earlier formulations in the case $x^{p^{r}}-a$ and the associated Wieferich‑type obstruction $a^{p}\equiv a\pmod{p^{2}}$.
A companion paper  [NGU25] by the first-named author studies the full $p$-adic shape of explicit integral bases for pure fields and proves that, for each $p^{e}\parallel n$, the local shape is determined by $a\bmod p^{e+1}$, yielding a minimal global period $M(n)=\prod_{p^{e}\parallel n}p^{e+1}=n\cdot\mathrm{rad}(n)$. In that framework the same local invariant $r_{p}(a)=\nu_{p}(a^{p-1}-1)-1$ governs denominators and gives a Wieferich-type threshold at $p\mid n$, which is exactly the quantity appearing in our $\alpha$-monogeneity criterion $\nu_{p}(m^{p}-m)=1$.
The density of square-free integers in progressions and basic $p$-adic features of Frobenius (Teichmüller lifts) will play a role in our distributional results, for accessible references see Brown’s paper [BRO21] and standard sources such as Cohen’s book [COH93] or Koblitz’s book [KOB84]. For pure cubics, monogeneity and quantitative aspects were studied by Aygin–Nguyen [AN21]. In the monic coefficient space, Bhargava–Shankar–Wang [BSW22, BSW25] proved that a positive density $\zeta(2)^{-1}$ of polynomials $f$ satisfy $\mathbb{Z}[x]/(f)=\mathcal{O}_{\mathbb{Q}[x]/(f)}$, and they determined the density of squarefree discriminants. Our work addresses instead the thin family $f_{m}(x)=x^{n}-m$: we give a complete local–global criterion for $\mathbb{Z}[\sqrt[n]{m}]$ to be maximal and compute the exact density $\frac{6}{\pi^{2}}\prod_{p\mid n}\frac{p}{p+1}$. Thus, Bhargava–Shankar–Wang treat the generic coefficient space (via invariant‑theoretic methods), while we handle a rigid one‑parameter family (via Dedekind’s criterion), and the two sets of results are complementary. From a moduli‑theoretic perspective, schemes parameterizing monogenic generators and notions of local monogenicity are developed in [ABS23], while not used here, this situates our local conditions in a geometric framework.
@@ 1.4. Acknowledgement
The first author thanks the Morningside Center of Mathematics, Chinese Academy of Sciences, for its support and a stimulating research environment. We thank Professor Győry for his interest in our work and for drawing our attention to his monograph with Evertse, as well as to the updated version of [EG24]. We also thank Professor Gaál for several helpful communications.
@@ 2. A criterion for $\alpha$-monogenic pure fields
For example, when $K=\mathbb{Q}(\alpha)$ with $\alpha\in\{\sqrt{-2},\sqrt[3]{3},\sqrt[5]{6}\}$, one checks that $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$.
Remark 2.3.
As Gaál kindly pointed out to us, the terminology $\alpha$-monogeneity is not standard in the literature. When a root $\alpha$ of a polynomial $f(x)$ generates a power integral basis, it is customary to say that the polynomial $f$ is monogenic, see, for example, the usage in Gassert’s work [GAS17]. We retain our terminology for clarity in the present context, but note this conventional viewpoint for the reader.
We record a version of Dedekind’s index theorem as follows.
@@ 3. Natural density for $\alpha$-monogenic pure fields
In this section we compute the natural density of integers $m$ for which $\mathcal{O}_{K}=\mathbb{Z}[\alpha]$ when $K$ is pure. Throughout, $n\geq 2$ is fixed.
@@
A set $E\subset\mathbb{Z}$ has natural density $\delta(E)$ if the following limit exists
$\delta(E):=\lim_{X\to\infty}\frac{1}{X}\#\{m\in E:\ 1\leq m\leq X\}.$
We write $\mu$ for the Möbius function, and $\mu^{2}$ for the indicator of square-free integers. We will also use the following elementary observations.
@@
Let $g(x)=x^{p}-x$. In $\mathbb{F}_{p}$ the roots are all $x\in\mathbb{F}_{p}$, at each root mod $p$ one has $g^{\prime}(x)=px^{p-1}-1\equiv-1\pmod{p}$, which is invertible. By Hensel’s lemma, each root lifts uniquely to $p^{k}$ for all $k$, so $\{x\bmod p^{2}:\ g(x)\equiv 0\}=p$. ∎
@@ 3.2. Global density via independence across primes
This is the key combinatorial identity underlying the main density theorem below. We also record that the restriction $x^{n}-m$ irreducible is asymptotically harmless as follows.
@@
Below are direct consequences of the density theorem.
@@ 3.3. Consequences and applications of the density formula
For questions beyond $\alpha$-monogeneity (e.g. tracking the full $p$-denominator profile of an integral basis), the canonical modulus is $M(n)=\prod_{p^{e}\parallel n}p^{e+1}=n\cdot\operatorname{rad}(n)$: the local $p$-shape depends only on $a\bmod p^{e+1}$ and this modulus is minimal. Our $\alpha$-monogeneity constraints at $p\mid n$ depend only on $a\bmod p^{2}$, but working modulo $M(n)$ aligns directly with the periodicity tables of the full shape, see [NGU25].